# DAVE: A VLM Vision Encoder for Document Understanding and Web Agents

### v2

@@ 1 Introduction
Vision-Language Models (VLMs) (liu2024llavanext; bai2025qwen2; achiam2023gpt) have shown remarkable capabilities in multi-modal reasoning and understanding, enabling a wide range of applications from image captioning to interactive web agents (wu2024atlas; qin2025ui; liu2023llava; li2022blip). A central component of VLMs is the vision encoder (yin2024survey), which converts images into visual tokens for the language backbone to process (li2023blip; liu2023llava). As such, the design of these encoders has become a key focus in VLM research (tong2024cambrian; shi2024eagle).
@@
However, the vision encoders predominantly used by VLMs suffer from a fundamental weakness: their low-level features lack the robust structural and spatial information essential for document understanding and web agents (tschannen2025siglip; radford2021learning). Conversely, DINO‑style models (oquab2023dinov2) offer low-level features yet are tuned to natural images (oquab2023dinov2; simeoni2025dinov3) and transfer poorly to documents, UIs, and diagrams (tong2024cambrian). To bridge this gap, we introduce DAVE (Document and web Agents Vision Encoder), a vision encoder purpose‑built for VLMs and tailored for these tasks. See Figure 1 for an overview.
@@
A key challenge in training on document and web images is the scarcity of high-quality annotated data, since current annotations rely on OCR models that bottleneck both scalability and quality (kim2021donut; lee2023pix2struct). We address this with a two-stage training process: self-supervised pretraining on large-scale unlabeled data, followed by autoregressive pretraining on limited annotated structural and grounding data. Yet, the supervised stage introduces a further challenge: specialized vision encoders trained with a single text decoder tend to overfit to that decoder, leading to misalignment when paired with other decoders. We mitigate this issue via multi-decoder pretraining and weight-merging (“model soup”) (wortsman2022model), which produces a largely decoder-agnostic encoder. To preserve high‑level semantics while leveraging our structural and spatial features, we fuse a generalist encoder with our encoder during pretraining via ensembling. The resulting specialized encoder integrates seamlessly with diverse decoders and generalist VLM stacks. Finally, our purpose-built vision encoder serves as the core component of a VLM architecture specifically designed for document and web agent tasks. We evaluate DAVE across diverse settings, including traditional and vision–language tasks. On traditional benchmarks, DAVE surpasses all SOTA models in document recognition and segmentation, while achieving competitive results with SigLIP 2 on screenshot classification. For vision–language evaluation, we test on document understanding, web localization, and web agent. Notably, DAVE improves performance by an average of 10.5% compared to SigLIP 2, highlighting its ability to provide specialized representations. Finally, we evaluate web agent settings on the Mind2Web benchmark (deng2023mind2web), where DAVE improves web agent performance by 5% over the strongest baseline vision encoder.
@@ 2 Related Work
Pretrained Vision Encoders. Image–text contrastive learning (radford2021learning; zhai2023sigmoid; jia2021scaling; yu2022coca) and self-supervised representation learning (he2022masked; zhou2021ibot; oquab2023dinov2; chen2020simple; caron2021emerging) are the two main paradigms for large-scale visual pretraining. Recent work has sought to combine the strengths of the two approaches through joint training or post-hoc alignment (tschannen2025siglip; fini2025multimodal; maninis2024tips; naeem2024silc), yielding strong encoders for both vision and multimodal tasks. The works most closely related to ours are Eagle  (shi2024eagle) and Perception Encoder  (bolya2025perception), which align pretrained vision encoders to a pretrained LLM with large-scale general datasets. In contrast, we pretrain the vision encoder from scratch with self-supervised learning, use multiple pretrained LLMs as a backbone, and use domain-specific data to create a specialized foundation model.
@@
Document Understanding with LLMs. Document processing and understanding have transitioned from OCR-based (xu2020layoutlm; xu2020layoutlmv2) methods to vision-language-based. One line of work trains encoder-decoder models with various objectives, like extracting structures (feng2025dolphin; kim2022ocr; lv2023kosmos; lee2023pix2struct), and image reconstruction conditioned on text (huang2022layoutlmv3; tang2023unifying). Recent advancements in VLMs led to another line of work that uses document data to finetune a VLM (nassar2025smoldocling; hu2024mplug; liu2024textmonkey) or incorporate it as part of the VLM training (team2025granite; bai2025qwen2; chen2024expanding). Our work seeks to bridge these two lines of work by introducing a specialized vision encoder within a general VLM framework.
@@
Vision-Language Models. Inspired by the success of recent large language models (LLMs) (brown2020language; dubey2024llama), Vision-language models (radford2021learning; liu2024llavanext; hua2024v2xum; ye2023mplug; tang2024avicuna; bai2025qwen2; zhu2025internvl3; hua2025finecaption; tong2024cambrian) aim to achieve multimodal intelligence by jointly understanding and generating visual and language information. Flamingo (alayrac2022flamingo) and BLIP-2 (li2023blip) are two of the early works that explore integrating LLMs as part of VLM. Beginning with LLaVA (liu2024llavanext), researchers have used instruction-following chat data in VQA format for instruction tuning, achieving significantly improved results (liu2024mmbench; li2023seed; hua2024mmcomposition; tang2025vidcomposition; lu2023mathvista).
@@
Multimodal Web Agents. Vision-based autonomous web agents have recently attracted attention for their simplicity and stronger generalization compared to LLM-based agents (deng2023mind2web; kim2023language). Early efforts such as WebGUM (furuta2023multimodal) and CogAgent (hong2024cogagent) pretrained vision–language models (VLMs) on web and GUI data to enhance agentic capabilities. Follow-up work (wu2024atlas; xu2024aguvis; liu2024harnessing; qin2025ui) expanded training with large-scale grounding and interaction datasets, further improving VLM performance on web-based tasks. While these works are promising steps toward a general-purpose visual web agent, the role of the vision encoder in these systems remains underexplored.
@@ 3.1 Preliminarilies
Masked Autoencoder. Masked Autoencoder (MAE) (he2022masked) adopts an asymmetric encoder–decoder architecture. The vision encoder processes only a subset of visible patches, while a lightweight transformer decoder reconstructs the full image from the encoded features and mask tokens. After pretraining, only the encoder is retained for downstream tasks.
@@ 3.2 Stage 1: SSL on Document and Web Images
In Stage 1, we conduct self-supervised training with MAE to learn rich structural and spatial features from document and web images. While MAE demonstrates the strongest performance on OCR tasks compared to other methods, the training is prone to instability when trained at scale (fan2025scaling). We find that this instability is particularly acute for document and web images. Our analysis (Section 5.4) attributes this to their characteristically low inter-patch variance, which destabilizes the target normalization in the standard MAE objective (Equation 1). To address this, we modify the objective to reconstruct raw pixel values directly:
@@ 4.1 Implementation Details
Self-supervised Stage. For the SSL stage, we follow the original MAE implementation (he2022masked), with ViT-L16-384 (dosovitskiy2020image) as our vision encoder. We train the MAE with a batch size of 4096 for 120K steps. More detail can be found in Appendix A.1
Supervised Training Stage. We adopt the VLM architecture discussed in Section 3.1, with an image tilting size of 384. During the ensemble training, we employ frozen SigLIP-2 as the generalist vision encoder, while the encoder from the Self-supervised stage serves as the domain-specialized component. This forms the full DAVE encoder. We experiment with multiple LLMs of varying scales and architectures, including QWen2.5-0.5B-Instruct (bai2025qwen2), Phi-4-mini-Instruct (abouelenin2025phi), and Granite-3.1-3B-Instruct (team2025granite). After a single epoch of full-parameter training, we retain only the DAVE encoder for weight-merging and downstream tasks. During the weight-merging distillation, we train the merge coefficient for 20 epochs on unlabeled documents and web images. Refer to Appendix B for implementation details.
@@ 4.2 Evaluation Setting
Baselines. We compare our vision encoder with both generalist and specialist encoders. For generalist encoders, we utilize DinoV2 (oquab2023dinov2) and Web-SSL MAE (fan2025scaling), a MAE variant trained on 2 billion images. We also include SigLIP2 (tschannen2025siglip) and AIMv2 (fini2025multimodal), which are SOTA encoders trained with both contrastive and reconstruction. For specialist vision encoders, we compare against DiT (li2022dit), Pix2Struct (lee2023pix2struct), and Dolphin (feng2025dolphin). Since both Pix2struct and Dolphin are encoder-decoder models, we use their encoders for comparison.
@@
Vision-Language Model as Evaluator. To evaluate our vision encoders on more complicated and realistic tasks, such as VQA and instruction-based localization, we use VLMs as evaluators. Specifically, we follow the standard LLaVA architecture as discussed in Section 3.1. We use a tilting size of 336 for AIMv2 and 384 for all other vision encoders. For vision encoders that produce a different number of visual tokens than the tilting resolution, we use bilinear interpolation as in LLaVA-Onevision (li2024llava) to interpolate the visual tokens. Note that we omit the projector alignment phase to improve training efficiency, as prior work has shown that this stage yields only limited gains tong2024cambrian; karamcheti2024prismatic. We use Llama-3.2-3B-Instruct (dubey2024llama) and Qwen2.5-7B-Instruct (bai2025qwen2) as the LLM backbone. We train for one epoch with the vision encoder frozen. For Mind2web, we finetune the VLMs on the training set before performing offline evaluation on the test set. More detail can be found in Appendix D.
@@ 4.3 Benchmarks
Classic Document Understanding Tasks. To directly evaluate the structural and spatial visual representation from the vision encoders, we consider several document parsing tasks that can be performed with only images. We employ DocBank (li2020docbank) to comprehensively evaluate document recognition on different categories, including tables, charts, and paragraphs. We also use Doclaynet (pfitzmann2022doclaynet) to test document segmentation. For web tasks, we use RICO-SCA (li2020mapping) to perform web UI classification.
Document and General VQA. For document understanding and question answering, we evaluate on AI2D (Kembhavi2016), OCRBench (Liu_2024), DocVQA (mathew2021docvqa), InfoVQA (Mathew_2022_WACV), and ChartQA (masry2022chartqa). For general vision-language understanding, we report results on MMMU (yue2024mmmu), RealWorldQA (xai2024realworldqa), and TextVQA (singh2019textvqa).
Web UI and Agent Tasks. For web UI localization, we use Screenspot-V2 (wu2024atlas), which consists of text instructions (e.g., “click the button in cooridnate (x, y)”), and the corresponding bounding boxes. We report the center accuracy: the label-box center lies within the predicted box. We also evaluate on WebSRC (chen2021websrc) and VisualWebBench (liu2024visualwebbench) for web UI question answering. For the agentic benchmark, we evaluate on Mind2Web (deng2023mind2web), where the model receives a user goal and a webpage screenshot, and must predict an action like clicking.
@@ 4.4 Datasets
Step SR Element Acc. Step SR Element Acc. Step SR Element Acc.
@@
Self-supervised Learning Data. We use 20 million images for our self-supervised training. For document data, we sample 10 million PDF images from DocFM (team2025granite), which contains 85 million document pages extracted from Common Crawl, Wikipedia, and ESG (Environmental, Social, and Governance) reports. The PDFs are filtered to include English only. For web screenshot, we sample 10 million from Common-Web (commonscreens) without any language filtering.
Supervised Learning Data. For the autoregressive supervised training, we leverage data from diverse sources, including PlotQA (methani2020plotqa), ChartQA (masry2022chartqa), fintabnet (zheng2021global), Datikz (belouadi2023automatikz), Pubtables (smock2022pubtables), and DocFM (team2025granite). Importantly, we use the version from Granite Vision (team2025granite) where all the problems are reformulated into content extraction. The data for this stage is a mixture of curated benchmarks and a large-scale, self-processed dataset. The curated benchmarks cover tasks such as chart-to-markdown, table-to-caption, and web UI grounding with UGround gou2024navigating. To expand on this, we processed 500K PDF samples from arXiv with an OCR model (cui2025paddleocr), formulating them into additional recognition and grounding tasks. Altogether, this training stage comprises approximately 2 million samples.
@@
Instruction Tuning Data. Following Cambrian-1 (tong2024cambrian), We use the LLaVA-1.5-mix 665K (liu2023llava), DocVQA (mathew2021docvqa), ChartQA (masry2022chartqa), and AI2D (Kembhavi2016) as our instruction tuning data. We also integrate Pixmo-Doc (yang2025scaling) into the data mix to improve document understanding. Additionally, we add MultiUI (liu2024harnessingwebpageuistextrich) to improve the web and agentic capability. This combined to a total of 2.5 million training images.

### v3

@@ Abstract
footnotetext: Equal advising. Correspondence to: zhaobin@berkeley.edu, hang.hua1@ibm.com
@@ 1 Introduction
Vision-Language Models (VLMs) (Liu et al., 2024a; Bai et al., 2025; Achiam et al., 2023) have shown remarkable capabilities in multi-modal reasoning and understanding, enabling a wide range of applications from image captioning to interactive web agents (Wu et al., 2024; Qin et al., 2025; Liu et al., 2023; Li et al., 2022b). A central component of VLMs is the vision encoder (Yin et al., 2024), which converts images into visual tokens for the language backbone to process (Li et al., 2023b; Liu et al., 2023). As such, the design of these encoders has become a key focus in VLM research (Tong et al., 2024; Shi et al., 2024).
@@
However, the vision encoders predominantly used by VLMs suffer from a fundamental weakness: their low-level features lack the robust structural and spatial information essential for document understanding and web agents (Tschannen et al., 2025; Radford et al., 2021). Conversely, DINO‑style models (Oquab et al., 2023) offer low-level features yet are tuned to natural images (Oquab et al., 2023; Siméoni et al., 2025) and transfer poorly to documents, UIs, and diagrams (Tong et al., 2024). To bridge this gap, we introduce DAVE (Document and web Agents Vision Encoder), a vision encoder purpose‑built for VLMs and tailored for these tasks. See Figure 1 for an overview.
@@
A key challenge in training on document and web images is the scarcity of high-quality annotated data, since current annotations rely on OCR models that bottleneck both scalability and quality (Kim et al., 2021; Lee et al., 2023). We address this with a two-stage training process: self-supervised pretraining on large-scale unlabeled data, followed by autoregressive pretraining on limited annotated structural and grounding data. Yet, the supervised stage introduces a further challenge: specialized vision encoders trained with a single text decoder tend to overfit to that decoder, leading to misalignment when paired with other decoders. We mitigate this issue via multi-decoder pretraining and weight-merging (“model soup”) (Wortsman et al., 2022), which produces a largely decoder-agnostic encoder. To preserve high‑level semantics while leveraging our structural and spatial features, we fuse a generalist encoder with our encoder during pretraining via ensembling. The resulting specialized encoder integrates seamlessly with diverse decoders and generalist VLM stacks. Finally, our purpose-built vision encoder serves as the core component of a VLM architecture specifically designed for document and web agent tasks. We evaluate DAVE across diverse settings, including traditional and vision–language tasks. On traditional benchmarks, DAVE surpasses all SOTA models in document recognition and segmentation, while achieving competitive results with SigLIP 2 on screenshot classification. For vision–language evaluation, we test on document understanding, web localization, and web agent. Notably, DAVE improves performance by an average of 10.5% compared to SigLIP 2, highlighting its ability to provide specialized representations. Finally, we evaluate web agent settings on the Mind2Web benchmark (Deng et al., 2023), where DAVE improves web agent performance by 5% over the strongest baseline vision encoder.
@@ 2 Related Work
Pretrained Vision Encoders. Image–text contrastive learning (Radford et al., 2021; Zhai et al., 2023; Jia et al., 2021; Yu et al., 2022) and self-supervised representation learning (He et al., 2022; Zhou et al., 2021; Oquab et al., 2023; Chen et al., 2020; Caron et al., 2021) are the two main paradigms for large-scale visual pretraining. Recent work has sought to combine the strengths of the two approaches through joint training or post-hoc alignment (Tschannen et al., 2025; Fini et al., 2025; Maninis et al., 2024; Naeem et al., 2024), yielding strong encoders for both vision and multimodal tasks. The works most closely related to ours are Eagle  (Shi et al., 2024) and Perception Encoder  (Bolya et al., 2025), which align pretrained vision encoders to a pretrained LLM with large-scale general datasets. In contrast, we pretrain the vision encoder from scratch with self-supervised learning, use multiple pretrained LLMs as a backbone, and use domain-specific data to create a specialized foundation model.
@@
Document Understanding with LLMs. Document processing and understanding have transitioned from OCR-based (Xu et al., 2020b; a) methods to vision-language-based. One line of work trains encoder-decoder models with various objectives, like extracting structures (Feng et al., 2025; Kim et al., 2022; Lv et al., 2023; Lee et al., 2023), and image reconstruction conditioned on text (Huang et al., 2022; Tang et al., 2023). Recent advancements in VLMs led to another line of work that uses document data to finetune a VLM (Nassar et al., 2025; Hu et al., 2024; Liu et al., 2024g) or incorporate it as part of the VLM training (Team et al., 2025; Bai et al., 2025; Chen et al., 2024). Our work seeks to bridge these two lines of work by introducing a specialized vision encoder within a general VLM framework.
@@
Vision-Language Models. Inspired by the success of recent large language models (LLMs) (Brown et al., 2020; Dubey et al., 2024), Vision-language models (Radford et al., 2021; Liu et al., 2024a; Sun et al., 2025; Hua et al., 2024a; Ye et al., 2023; Tang et al., 2024; Bai et al., 2025; Zhu et al., 2025; Hua et al., 2025; Tong et al., 2024) aim to achieve multimodal intelligence by jointly understanding and generating visual and language information. Flamingo (Alayrac et al., 2022) and BLIP-2 (Li et al., 2023b) are two of the early works that explore integrating LLMs as part of VLM. Beginning with LLaVA (Liu et al., 2024a), researchers have used instruction-following chat data in VQA format for instruction tuning, achieving significantly improved results (Liu et al., 2024e; Li et al., 2023a; Hua et al., 2024b; Tang et al., 2025; Lu et al., 2023).
@@
Multimodal Web Agents. Vision-based autonomous web agents have recently attracted attention for their simplicity and stronger generalization compared to LLM-based agents (Deng et al., 2023; Kim et al., 2023). Early efforts such as WebGUM (Furuta et al., 2023) and CogAgent (Hong et al., 2024) pretrained vision–language models (VLMs) on web and GUI data to enhance agentic capabilities. Follow-up work (Wu et al., 2024; Xu et al., 2024; Liu et al., 2024b; Qin et al., 2025) expanded training with large-scale grounding and interaction datasets, further improving VLM performance on web-based tasks. While these works are promising steps toward a general-purpose visual web agent, the role of the vision encoder in these systems remains underexplored.
@@ 3.1 Preliminarilies
Masked Autoencoder. Masked Autoencoder (MAE) (He et al., 2022) adopts an asymmetric encoder–decoder architecture. The vision encoder processes only a subset of visible patches, while a lightweight transformer decoder reconstructs the full image from the encoded features and mask tokens. After pretraining, only the encoder is retained for downstream tasks.
@@ 3.2 Stage 1: SSL on Document and Web Images
In Stage 1, we conduct self-supervised training with MAE to learn rich structural and spatial features from document and web images. While MAE demonstrates the strongest performance on OCR tasks compared to other methods, the training is prone to instability when trained at scale (Fan et al., 2025). We find that this instability is particularly acute for document and web images. Our analysis (Section 5.4) attributes this to their characteristically low inter-patch variance, which destabilizes the target normalization in the standard MAE objective (Equation 1). To address this, we modify the objective to reconstruct raw pixel values directly:
@@ 4.1 Implementation Details
Self-supervised Stage. For the SSL stage, we follow the original MAE implementation (He et al., 2022), with ViT-L16-384 (Dosovitskiy et al., 2020) as our vision encoder. We train the MAE with a batch size of 4096 for 120K steps. More detail can be found in Appendix A.1
Supervised Training Stage. We adopt the VLM architecture discussed in Section 3.1, with an image tilting size of 384. During the ensemble training, we employ frozen SigLIP-2 as the generalist vision encoder, while the encoder from the Self-supervised stage serves as the domain-specialized component. This forms the full DAVE encoder. We experiment with multiple LLMs of varying scales and architectures, including QWen2.5-0.5B-Instruct (Bai et al., 2025), Phi-4-mini-Instruct (Abouelenin et al., 2025), and Granite-3.1-3B-Instruct (Team et al., 2025). After a single epoch of full-parameter training, we retain only the DAVE encoder for weight-merging and downstream tasks. During the weight-merging distillation, we train the merge coefficient for 20 epochs on unlabeled documents and web images. Refer to Appendix B for implementation details.
@@ 4.2 Evaluation Setting
Baselines. We compare our vision encoder with both generalist and specialist encoders. For generalist encoders, we utilize DinoV2 (Oquab et al., 2023) and Web-SSL MAE (Fan et al., 2025), a MAE variant trained on 2 billion images. We also include SigLIP2 (Tschannen et al., 2025) and AIMv2 (Fini et al., 2025), which are SOTA encoders trained with both contrastive and reconstruction. For specialist vision encoders, we compare against DiT (Li et al., 2022a), Pix2Struct (Lee et al., 2023), and Dolphin (Feng et al., 2025). Since both Pix2struct and Dolphin are encoder-decoder models, we use their encoders for comparison.
@@
Vision-Language Model as Evaluator. To evaluate our vision encoders on more complicated and realistic tasks, such as VQA and instruction-based localization, we use VLMs as evaluators. Specifically, we follow the standard LLaVA architecture as discussed in Section 3.1. We use a tilting size of 336 for AIMv2 and 384 for all other vision encoders. For vision encoders that produce a different number of visual tokens than the tilting resolution, we use bilinear interpolation as in LLaVA-Onevision (Li et al., 2024) to interpolate the visual tokens. Note that we omit the projector alignment phase to improve training efficiency, as prior work has shown that this stage yields only limited gains Tong et al. (2024); Karamcheti et al. (2024). We use Llama-3.2-3B-Instruct (Dubey et al., 2024) and Qwen2.5-7B-Instruct (Bai et al., 2025) as the LLM backbone. We train for one epoch with the vision encoder frozen. For Mind2web, we finetune the VLMs on the training set before performing offline evaluation on the test set. More detail can be found in Appendix D.
@@ 4.3 Benchmarks
Classic Document Understanding Tasks. To directly evaluate the structural and spatial visual representation from the vision encoders, we consider several document parsing tasks that can be performed with only images. We employ DocBank (Li et al., 2020a) to comprehensively evaluate document recognition on different categories, including tables, charts, and paragraphs. We also use Doclaynet (Pfitzmann et al., 2022) to test document segmentation. For web tasks, we use RICO-SCA (Li et al., 2020b) to perform web UI classification.
Document and General VQA. For document understanding and question answering, we evaluate on AI2D (Kembhavi et al., 2016), OCRBench (Liu et al., 2024f), DocVQA (Mathew et al., 2021), InfoVQA (Mathew et al., 2022), and ChartQA (Masry et al., 2022). For general vision-language understanding, we report results on MMMU (Yue et al., 2024), RealWorldQA (xAI, 2024), and TextVQA (Singh et al., 2019).
Web UI and Agent Tasks. For web UI localization, we use Screenspot-V2 (Wu et al., 2024), which consists of text instructions (e.g., “click the button in cooridnate (x, y)”), and the corresponding bounding boxes. We report the center accuracy: the label-box center lies within the predicted box. We also evaluate on WebSRC (Chen et al., 2021) and VisualWebBench (Liu et al., 2024d) for web UI question answering. For the agentic benchmark, we evaluate on Mind2Web (Deng et al., 2023), where the model receives a user goal and a webpage screenshot, and must predict an action like clicking.
@@ 4.4 Datasets
Element Acc. Step SR Element Acc. Step SR Element Acc. Step SR
@@
Self-supervised Learning Data. We use 20 million images for our self-supervised training. For document data, we sample 10 million PDF images from DocFM (Team et al., 2025), which contains 85 million document pages extracted from Common Crawl, Wikipedia, and ESG (Environmental, Social, and Governance) reports. The PDFs are filtered to include English only. For web screenshot, we sample 10 million from Common-Web (com, 2023) without any language filtering.
Supervised Learning Data. For the autoregressive supervised training, we leverage data from diverse sources, including PlotQA (Methani et al., 2020), ChartQA (Masry et al., 2022), fintabnet (Zheng et al., 2021), Datikz (Belouadi et al., 2023), Pubtables (Smock et al., 2022), and DocFM (Team et al., 2025). Importantly, we use the version from Granite Vision (Team et al., 2025) where all the problems are reformulated into content extraction. The data for this stage is a mixture of curated benchmarks and a large-scale, self-processed dataset. The curated benchmarks cover tasks such as chart-to-markdown, table-to-caption, and web UI grounding with UGround Gou et al. (2024). To expand on this, we processed 500K PDF samples from arXiv with an OCR model (Cui et al., 2025), formulating them into additional recognition and grounding tasks. Altogether, this training stage comprises approximately 2 million samples.
@@
Instruction Tuning Data. Following Cambrian-1 (Tong et al., 2024), We use the LLaVA-1.5-mix 665K (Liu et al., 2023), DocVQA (Mathew et al., 2021), ChartQA (Masry et al., 2022), and AI2D (Kembhavi et al., 2016) as our instruction tuning data. We also integrate Pixmo-Doc (Yang et al., 2025) into the data mix to improve document understanding. Additionally, we add MultiUI (Liu et al., 2024c) to improve the web and agentic capability. This combined to a total of 2.5 million training images.